Harry pushes his 1.50x103 N school trunk with a horizontal force of 345 N for a distance of 24.0 m through
the train station. The coefficient of kinetic friction between the trunk and the ground is 0.22. Using work and
energy techniques, find the speed of the trunk after moving 24.0 m, assuming the trunk starts at rest.
Answer #1

The resisting force of fiction = mu*N

N is normal reaction, here it is trunks weight itself

m = 1500/g = 150 kg

f = mu*mg = 0.22*1.5e3 = 330 N

applied force F = 345 N

F - f = ma

a = (345-330)/15 = 1m/s^2

using energy and work principle

work done by net force results in kinetic energy of trunk

ma*d = 1/2 mv^2

v = sqrt(2ad) = 6.93 m/s

Please upvote if you like the answer!!